OpenStudy (anonymous):

One kilogram is about 2.205 pounds.when a 55 gallon aquarium is filled it weighs 625 pounds. To the NEAREST TENTH how many kilograms is this.?

OpenStudy (anonymous):

So just divide 625 by 2.205 = 283.5 kg

OpenStudy (anonymous):

To do this, use elimination. \[\frac{ 1kilogram }{ 2.205pounds } = \frac{ 625pounds }{ 55gallons }\] multiplying this out cancels out the pounds and leaves you with \[\frac{ 283.45kilograms }{ 55gallons }\] So you're answer is in a 55 gallon tank, it weighs 283.45kilograms.

OpenStudy (anonymous):

Thanks, alot :D